Please help.
A cockroach can run at about 2mph. The average body length of a cockroach is about 1 inch. The average body length of a horse is 8 feet. If we asume that a horse's speed-to-length ratio is the same as that of a cockroach, how fast can a horse run?

Answer by sarah_adam(201)

cockroach = speed : length = 2:1 ( 1 in )
horse = speed : length = s:8 (8ft )
so lets convert ft to in
1ft = 12 in
8ft = 8 *12 = 96in
now according to the assumption
2:1 = s:96
2/1 = s/96
s = 2*96= 192 mph
speed of horse = 192 mph